OpenStudy (anonymous):

how do I simplify the following: (a^-5b)^-4/ab^3

OpenStudy (amistre64):

thats hard to decipher into a problem... can you re write it better or explain it more?

OpenStudy (amistre64):

is this it? \[\frac{(a^{-5b})^{-4}}{ab^3}\]

OpenStudy (anonymous):

the slash after the 4 is actually indicative of the fraction. wasn't sure how to write that

OpenStudy (anonymous):

yes that is it

OpenStudy (amistre64):

exponents of exponents multiply together... a^(20b) ------- ab^3

OpenStudy (amistre64):

division of like bases subtrats exponents: a^b --- = a^(b-c) a^c

OpenStudy (amistre64):

a^(20b-1) --------- seems to be the best I can get from it... b^3

OpenStudy (amistre64):

what constitutes a simplification ?

OpenStudy (anonymous):

I got a^20 over ab^-12 but not sure if I did it right

OpenStudy (amistre64):

is there a choice of answers to base this off of?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its hard to write equation on here lol

OpenStudy (amistre64):

we could try log rules.... maybe :) 20b log(a) - log(a) + 3log(b)...not sure how that would help tho :)

OpenStudy (amistre64):

the equation editor on the bottom left helps alittle

OpenStudy (anonymous):

its my first time using this site

OpenStudy (amistre64):

takes a little to get used to but its handy at times

OpenStudy (amistre64):

it slows down the computer tho... lol

OpenStudy (anonymous):

I noticed. I tried to use the equation template but it got confusing

OpenStudy (amistre64):

you can make a fraction by typing 'frac{top}{bottom}' into it directly

OpenStudy (amistre64):

exponents are ^{number}

OpenStudy (anonymous):

\[(a^-5b)^-4 \ab^3\]

OpenStudy (anonymous):

oops that didn't come out right

OpenStudy (amistre64):

sooo close.....

OpenStudy (amistre64):

type it in like like this: frac{a^{-5b}^{-4}}{ab^3}

OpenStudy (anonymous):

I have to go...phone

OpenStudy (anonymous):

thanks for your help :)